# Kaiwu: A Multimodal Manipulation Dataset and Framework for Robot Learning and Human-Robot Interaction

### v1

@@ VI-B2 Glove data
For the assembly actions of C1-C15, the data gloves are transferred and stored as r/l_TIME.csv files with columns 1-12 representing the quaternion data of the palm, forearm, and upper arm, columns 13-31 representing the finger angle sensor data, and columns 32-50 representing the grip force sensor values. In addition, the data can be exported to present a visualization interface of the hand activity during data collection, and stored as an MP4 file. The GloveData directory structure is shown in Fig. 12.
{forest}
forest scheme [glove data [glove_data [left [l_TIME.csv (C1)] [⋮] [l_TIME.csv (C15)] ] [right [r_TIME.csv (C1)] [⋮] [r_TIME.csv (C15)] ] ] [glove_export [C1.mp4] [⋮] [C15.mp4] ] ]
Figure 12: Structure of the glove data.
@@ VI-B3 RGB-D data
RGB-D data is stored in the C1-C15.mkv files in the RGB_video file and contains the timestamps used to generate the picture data, recorded as .txt files. The data generated based on the timestamp is stored in the IMG_picture file, including RGB images (.jpg file and .pcd file), depth information (.png file). The environmental information of the assembly data of C1-C15 is stored under 15 folders respectively. The kinect_data directory structure is shown in Fig. 13.
{forest}
forest scheme [RGB-D data [IMG_picture [C1 [color.jpg] [depth.png] [pcd] ] [⋮] [C15_source_data]] [RGB_video [C1.mkv] [C1.txt] [⋮] [C15.mkv] [C15.txt] ] ]
Figure 13: Structure of the RGB-D data.
@@ VI-B4 Ground truth data
After the initial body joint point calibration, the NOKOV system captures and records the 3D spatial coordinates of each marker point as the assembly process proceeds, outputting the data as a (.cap .trb .xrb) File. Visualization is achieved by reimporting the data. Ground truth data directory is structured as Fig. 14.
{forest}
forest scheme [skeleton data [BodyNUM.mars] [Remotetake1.cap] [Remotetake1-Unnamed.trb] [Remotetake1-BodyNUM.trb] [Remotetake1-BodyNUM.xrb] [⋮] [Remotetake15.cap] [Remotetake15-Unnamed.trb] [Remotetake15-BodyNUM.trb] [Remotetake15-BodyNUM.xrb] ]
Figure 14: Structure of the skeletons data.
@@ VI-B5 Eye tracker data
The Tobii-data folder contains the first-person perspective information (tobii-data) and the eye-tracking information (tobii-export). The former contains the acquisition data from C1 to C15 and generates the corresponding .gz data compression files (eventdata records the events, gazedata records the eye-tracking sensor data, and imudata records the IMU measurement data, which are combined and displayed in the tobi-export file) and the visualization .mp4 file. The latter contains C1-C15 .mp4 with gaze point movements, and .xlsx files listing the data tables measured by the built-in gyroscope, magnetometer, accelerometer and eye tracker. Each header in the files is labeled with the corresponding data meaning. The directory structure is shown in Fig. 15.
{forest}
forest scheme [eye
tracker
data [tobii_data [C1_source_data [eventdata.gz] [gazedata.gz] [imudata.gz] [recording.g3] [scenevideo.mp4] ] [⋮] [C15_source_data] ] [tobii_export [C1 [C1.mp4] [Data Export.xlsx] ] [⋮] [C15 [C15.mp4] [Data Export.xlsx] ] ] ]
Figure 15: Structure of the eye tracker data.
@@ VI-B6 Voice data
The microphone outputs environment sound information as .wav audio files and .txt timestamp data. Microphones numbered 1-2 record the ambient sound in the assembly area and 3-4 record the ambient sound in the accessory area. The audio data directory structure is shown in Fig. 16.
{forest}
forest scheme [voice data [micophione1 [C1_TIME.wav] [C1_TIME.txt] [⋮] [C15_TIME.wav] [C15_TIME.txt] ] [⋮] [micophione4 [C1_TIME.wav] [C1_TIME.txt] [⋮] [C15_TIME.wav] [C15_TIME.txt] ] ]
Figure 16: Structure of the vioce data.
@@ VII Accessing the data
The data has been uploaded to ScienceDB with DOI number https://doi.org/10.57760/sciencedb.14937. The homepage of this dataset can be searched using the keyword ”Kaiwu” on ScienceDB. The following files are provided for download: A compilation of all the data Kaiwu Data Annotation (31.84GB), which action segmentation (963MB), contains the timestamp exported delsys_data with ACC and EMG and the gloves_data with action labels. The AOIs (17.9GB) contains the annotation results and annotation process files which that can be opened using Tobii Lab software. The gesture segmentation (973MB) contains the delsys_data and gloves_data with fine grain gesture labels. The semantic segmentation (11.6GB) contains the part1-4.zip with RGB video annotation results, and finally gesture classification (417MB), contains the hand state classification data(classification result.zip).
@@
VIII Known issues
Owing to limitation of computational load, certain data output streams may have periodic interruptions. The coverage of available multi-modal data is recorded in the descriptive statistics (Table VI), to assess the percentage of data stream dropouts.
Since the data gloves are not equipped with sensors at the fingertips, although participants are informed to minimize the use of bare fingertip areas not equipped with sensors, a small amount of this phenomenon still occurred in the experiments, resulting in incomplete data collection for a small portion of the glove sensing data.
IX Conclusion

### v2

@@ VI-B2 Glove data
For the assembly actions of C1-C15, the data gloves are transferred and stored as r/l_TIME.csv files with columns 1-12 representing the quaternion data of the palm, forearm, and upper arm, columns 13-31 representing the finger angle sensor data, and columns 32-50 representing the grip force sensor values. In addition, the data can be exported to present a visualization interface of the hand activity during data collection, and stored as an MP4 file.
@@ VI-B3 RGB-D data
RGB-D data is stored in the C1-C15.mkv files in the RGB_video file and contains the timestamps used to generate the picture data, recorded as .txt files. The data generated based on the timestamp is stored in the IMG_picture file, including RGB images (.jpg file and .pcd file), depth information (.png file). The environmental information of the assembly data of C1-C15 is stored under 15 folders respectively.
@@ VI-B4 Ground truth data
After the initial body joint point calibration, the Motion Capture system captures and records the 3D spatial coordinates of each marker point as the assembly process proceeds, outputting the data as a (.cap .trb .xrb) File. Visualization is achieved by reimporting the data.
@@ VI-B5 Eye tracker data
The Tobii-data folder includes first-person perspective data (tobii-data) and eye-tracking data (tobii-export). Tobii-data contains compressed .gz files (eventdata for events, gazedata for eye-tracking, imudata for IMU measurements) and a visualization .mp4 file. Tobii-export includes .mp4 files showing gaze point movements for C1-C15 and .xlsx files with data tables from the gyroscope, magnetometer, accelerometer, and eye tracker. Each file header is labeled with the corresponding data meaning.
@@ VI-B6 Voice data
The microphone outputs environment sound information as .wav audio files and .txt timestamp data. Microphones numbered 1-2 record the ambient sound in the assembly area and 3-4 record the ambient sound in the accessory area.
@@ VII Accessing the data
The data has been uploaded to ScienceDB with DOI number. The homepage of this dataset can be searched using the keyword ”Kaiwu” on ScienceDB. The following files are provided for download: A compilation of all the data Kaiwu Data Annotation (31.84GB), which action segmentation (963MB), contains the timestamp exported delsys_data with ACC and EMG and the gloves_data with action labels. The AOIs (17.9GB) contains the annotation results and annotation process files which that can be opened using Tobii Lab software. The gesture segmentation (973MB) contains the delsys_data and gloves_data with fine grain gesture labels. The semantic segmentation (11.6GB) contains the part1-4.zip with RGB video annotation results, and finally gesture classification (417MB), contains the hand state classification data(classification result.zip).
@@
VIII Conclusion